what is the final pressure of 1.5 liters of an ideal gas if the initial volume at 5.0 atmosphere was 3.0L?
1
Expert's answer
2019-05-23T07:58:27-0400

pV = const

p1= 5 atm

V2 = 1.5 l

V1 = 3 atm

p2 = (p1V1)/V2 = (5*3)/1.5 = 10 atm